# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ A.1 Empirical Wasserstein-2 convergence
We will use below the following fact. See [32, Appendix E] and references within for its justification.
@@
The following results are from [32, Appendix E].
@@ A.2 Proximal map
This follows from the assumption that $h(x)$ is twice continuously differentiable on $\mathfrak{D}^{c}$ and the implicit differentiation calculation shown in [40, Appendix B1]. For $x\in\{x:\operatorname{Prox}_{vh}(x)\in\mathfrak{D}\}$, $\operatorname{Prox}_{vh}(x)$ is differentiable and has derivative equal to 0 except for a finite set of points. To see this, note that preimage $\operatorname{Prox}^{-1}_{vh}(\mathbf{y})$ for $y\in\mathfrak{D}$ is either a singleton set or a closed interval of the form $[x_{1},x_{2}]$ for $x_{1}\in\mathbb{R}\cup\{-\infty\},x_{2}\in\mathbb{R}\cup\{+\infty\}$ and $x_{1}<x_{2}$, using continuity and monotonicity of $x\mapsto\operatorname{Prox}_{vh}(x)$. This implies that $\{x:\operatorname{Prox}_{vh}(x)\in\mathfrak{D}\}$ is a union of finite number of singleton sets and a finite number of closed intervals. Furthermore, $\operatorname{Prox}_{vh}(x)$ is constant on each of the closed intervals. It follows that $\operatorname{Prox}_{vh}(x)$ is differentiable and has derivative equal to 00 on the interiors of the closed intervals, and that $\mathcal{C}$ is union of some of the singleton sets and all of the finite-valued endpoints of the closed intervals.
@@
Note that we have assigned +$\infty$+ ∞ to $h^{\prime\prime}$ on $\mathfrak{D}$ and 00 to $\operatorname{Prox}^{\prime}_{vh}$ on $\mathcal{C}$. Piecewise continuity of $x\mapsto\frac{1}{w+h^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}$ for any $w>0$ follows from the discussion above. ∎
@@ A.3 Properties of R- and Cauchy transform
The following shows that the Cauchy- and R-transforms of $-\mathsf{D}^{2}$ are well-defined by (14), and reviews their properties.
@@
Let $G(\cdot)$ and $R(\cdot)$ be the Cauchy- and R-transforms of $-\mathsf{D}^{2}$ under Assumption 2.
@@
    The function $G:(-d_{-},\infty)\to\mathbb{R}$ is positive and strictly decreasing. Setting $G\left(-d_{-}\right):=\lim_{z\to-d_{-}}G(z)\in(0,\infty]$, $G$ admits a functional inverse $G^{-1}:(0,G(-d_{-}))\to(-d_{-},\infty)$.
@@
    where $\left\{\kappa_{k}\right\}_{k\geq 1}$ are the free cumulants of the law of $-\mathsf{D}^{2}$ and $\kappa_{1}=-\mathbb{E}\mathsf{D}^{2}$ and $\kappa_{2}=\mathbb{V}(\mathsf{D}^{2})$.
@@
See [59, Lemma G.6] for (a) and (b), To see (c), for any $z\in\left(0,G\left(-d_{-}\right)\right)$, differentiating $R(z)=G^{-1}(z)-z^{-1}$ yields
$-zR^{\prime}(z)=z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)^{-1}-\frac{1}{z}$
@@
$\Leftrightarrow z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)>G^{-1}(z)$
@@
where we used in the second line that $R(z)=G^{-1}(z)-1/z<0$ from (b). Note that the last line is true since $\mathsf{D}^{2}\neq 0$ with positive probability. (e) trivially follows from (c). (f) follows from [69, Notation 12.6, Proposition 13.15]. ∎
@@ A.4 DF adjustment coincide with Spectrum-Aware adjustment under Marchenko-Pastur law
If the empirical distribution of the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$ weakly converges Marchenko-Pastur law, then |$\absolutevalue{\widehat{\mathsf{adj}}-\breve{\mathsf{adj}}}\to 0.$
@@
$01+\left(\lambda_{2}+1-\delta^{-1}\right)G\left(-\lambda_{2}\right)-\delta^{-1}\lambda_{2}\quantity(G\left(-\lambda_{2}\right))^{2}=$ 0 . (48)

### v6

@@ A.1 Empirical Wasserstein-2 convergence
We will use below the following fact. See [39, Appendix E] and references within for its justification.
@@
The following results are from [39, Appendix E].
@@ A.2 Proximal map
  • (c)
    We have the following inequality
    $\left|\operatorname{Prox}_{v_{1}h}(x)-\operatorname{Prox}_{v_{2}h}(x)\right|\leq\frac{\left|v_{1}-v_{2}\right|}{\left(1+v_{1}c_{0}\right)v_{2}}\left|x-\operatorname{Prox}_{v_{2}h}(x)\right|.$
  • (d)
    For any $v>0$, it cannot be true that $\operatorname{Prox}_{vh}^{\prime}(x)=1$ almost everywhere.
Proof of Proposition A.7.
We skip proofs of (a) and (b) which are well-known properties of proximal operator. We now prove (c). Let $y_{1}=\operatorname{Prox}_{v_{1}h}(x)$ and $y_{2}=\operatorname{Prox}_{v_{2}h}(x)$. By the definition of the proximal mapping, there exist subgradients $u_{1}\in\partial h(y_{1})$ and $u_{2}\in\partial h(y_{2})$ such that
$x-y_{1}=v_{1}\,u_{1}\quad\text{and}\quad x-y_{2}=v_{2}\,u_{2}.$
Subtracting these two relations yields
$y_{2}-y_{1}=v_{1}(u_{1}-u_{2})+(v_{1}-v_{2})u_{2}.$
Since we are in the scalar setting, we can multiply both sides by $y_{1}-y_{2}$ (note that $(y_{2}-y_{1})(y_{1}-y_{2})=-|y_{1}-y_{2}|^{2}$) to obtain
$|y_{1}-y_{2}|^{2}=v_{1}(u_{1}-u_{2})(y_{1}-y_{2})+(v_{1}-v_{2})u_{2}(y_{1}-y_{2}).$
The convexity assumption on $h$ with parameter $c_{0}\leq 0$ implies that for any $u_{1}\in\partial h(y_{1})$ and $u_{2}\in\partial h(y_{2})$ we have
$(u_{1}-u_{2})(y_{1}-y_{2})\geq c_{0}\,(y_{1}-y_{2})^{2}.$
Substituting this inequality into the previous display yields
$|y_{1}-y_{2}|^{2}\geq v_{1}\,c_{0}\,(y_{1}-y_{2})^{2}+(v_{1}-v_{2})u_{2}(y_{1}-y_{2}).$
Rearranging terms and taking absolute values, we deduce
$(1+v_{1}c_{0})|y_{1}-y_{2}|^{2}\leq|v_{1}-v_{2}|\;|u_{2}|\;|y_{1}-y_{2}|.$
Assuming $y_{1}\neq y_{2}$ so that we can cancel a factor of $|y_{1}-y_{2}|$, it follows that
$|y_{1}-y_{2}|\leq\frac{|v_{1}-v_{2}|}{1+v_{1}c_{0}}\,|u_{2}|.$
Finally, recalling that the optimality condition for $y_{2}$ gives $x-y_{2}=v_{2}\,u_{2}$, we have $|u_{2}|=\frac{|x-y_{2}|}{v_{2}}$. Substituting this expression into the inequality above yields
$|y_{1}-y_{2}|\leq\frac{|v_{1}-v_{2}|}{(1+v_{1}c_{0})v_{2}}\,|x-y_{2}|,$
which is the desired result.
To see (d), note that since Assumption 3 requires $h(x)$ to be non-constant, it suffices to show the following: If $\operatorname{Prox}^{\prime}_{vh}(x)=1$ almost everywhere $v>0$, then $h(x)$ is constant. Define
$T(x)=\operatorname{Prox}_{vh}(x).$
Since $T^{\prime}(x)=1$ almost everywhere, integrating over an interval shows that
$T(b)-T(a)=b-a\quad\text{for all }a,b\in\mathbb{R}.$
Thus, $T$ is an affine function of the form
$T(x)=x+c,$
for some constant $c\in\mathbb{R}$.
Now, by the definition of the proximal operator, for every $x\in\mathbb{R}$ the optimality condition (in terms of subgradients) implies that
$0\in\partial h\bigl(T(x)\bigr)+\frac{1}{v}\bigl(T(x)-x\bigr).$
Substituting $T(x)=x+c$ gives
$0\in\partial h(x+c)+\frac{c}{v},\quad\text{for all }x\in\mathbb{R}.$
Letting $y=x+c$, we deduce that for every $y\in\mathbb{R}$
$\frac{c}{v}\in\partial h(y).$
By our assumption, $h$ is twice continuously differentiable except for a finite set of points. Integrating, we obtain
$h(y)=-\frac{c}{v}y+b,$
for some constant $b\in\mathbb{R}$. However, $h$ is assumed to be nonnegative. Hence, we must have $c=0$.
@@
This follows from the assumption that $h(x)$ is twice continuously differentiable on $\mathfrak{D}^{c}$ and the implicit differentiation calculation shown in [48, Appendix B1]. For $x\in\{x:\operatorname{Prox}_{vh}(x)\in\mathfrak{D}\}$, $\operatorname{Prox}_{vh}(x)$ is differentiable and has derivative equal to 0 except for a finite set of points. To see this, note that preimage $\operatorname{Prox}^{-1}_{vh}(\mathbf{y})$ for $y\in\mathfrak{D}$ is either a singleton set or a closed interval of the form $x_{1},x_{2}]$ for $x_{1}\in\mathbb{R}\cup\{-\infty\},x_{2}\in\mathbb{R}\cup\{+\infty\}$ and $x_{1}<x_{2}$, using continuity and monotonicity of $x\mapsto\operatorname{Prox}_{vh}(x)$. This implies that $\{x:\operatorname{Prox}_{vh}(x)\in\mathfrak{D}\}$ is a union of finite number of singleton sets and a finite number of closed intervals. Furthermore, $\operatorname{Prox}_{vh}(x)$ is constant on each of the closed intervals. It follows that $\operatorname{Prox}_{vh}(x)$ is differentiable and has derivative equal to 0 on the interiors of the closed intervals, and that $\mathcal{C}$ is union of some of the singleton sets and all of the finite-valued endpoints of the closed intervals.
@@
Note that we have assigned $\infty$ to $h^{\prime\prime}$ on $\mathfrak{D}$ and 0 to $\operatorname{Prox}^{\prime}_{vh}$ on $\mathcal{C}$. Piecewise continuity of $x\mapsto\frac{1}{w+h^{\prime\prime}\left(\operatorname{Prox}_{vh}(x)\right)}$ for any $w>0$ follows from the discussion above. ∎
@@ A.3 Properties of R- and Cauchy transform
The following shows that the Cauchy- and R-transforms of $\mathsf{D}^{2}$ are well-defined by (14), and reviews their properties.
@@
Let $G(\cdot)$ and $R(\cdot)$ be the Cauchy- and R-transforms of $\mathsf{D}^{2}$ under Assumption 1.
@@
    The function $G:(-d_{-},\infty)\to\mathbb{R}$ is positive and strictly decreasing. Setting $\left(-d_{-}\right):=\lim_{z\to-d_{-}}G(z)\in(0,\infty], GG$ admits a functional inverse $G^{-1}:(0,G(-d_{-}))\to(-d_{-},\infty)$.
@@
    where $\left\{\kappa_{k}\right\}_{k\geq 1}$ are the free cumulants of the law of $\mathsf{D}^{2}$ and $\kappa_{1}=-\mathbb{E}\mathsf{D}^{2}$ and $\kappa_{2}=\mathbb{V}(\mathsf{D}^{2})$.
@@
See [69, Lemma G.6] for (a) and (b), To see (c), for any $z\in\left(0,G\left(-d_{-}\right)\right)$, differentiating $R(z)=G^{-1}(z)-z^{-1}$ yields
$zR^{\prime}(z)=z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)^{-1}-\frac{1}{z}$
@@
$\Leftrightarrow z\left(\mathbb{E}\frac{1}{\left(\mathsf{D}^{2}+G^{-1}(z)\right)^{2}}\right)^{-1}>G^{-1}(z)$
@@
where we used in the second line that $R(z)=G^{-1}(z)-1/z<0$ from (b). Note that the last line is true since $\mathsf{D}^{2}\neq 0$ with positive probability. (e) trivially follows from (c). (f) follows from [84, Notation 12.6, Proposition 13.15]. ∎
@@ A.4 DF adjustment coincide with Spectrum-Aware adjustment under Marchenko-Pastur law
[…]
@@
$1+\left(\lambda_{2}+1-\delta^{-1}\right)G\left(-\lambda_{2}\right)-\delta^{-1}\lambda_{2}\quantity(G\left(-\lambda_{2}\right))^{2}=0.$ (39)